What volume (in mL) of 8.84 M HBr would be required to make 300.01 mL of a solution with a pH of 2.57?
Answer #1

pH = 2.57, PH=-149CHT) = 2.57 * CH+] = 7-69x150 mm - HBr - Ht tor- = CH Br] = R-698163m - Vi=? , Mi=8-84 m HBV dilution - V2